# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ C.2 Proof of Lemma 8
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}\left(\psi_{\bm{u}/\sqrt{N}}^{\otimes N}\right),\phi_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=0$ (296)
for $\epsilon\in(0,1/6)$, where $\psi_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}\psi e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ and $\phi_{\bm{\theta}}\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}\phi e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}$ for $\bm{\theta}\in\mathbb{R}^{m}$.
Proof.
Following Lemma 30, we introduce $\tilde{X}_{i}\coloneqq(I-\psi)X_{i}\psi+\psi X_{i}(I-\psi)$ and $\tilde{X}_{i}^{\prime}\coloneqq(I-\phi)X_{i}^{\prime}\phi+\phi X_{i}^{\prime}(I-\phi)$ and denote
$\tilde{\psi}_{\bm{u}/\sqrt{N}}\coloneqq e^{\mathrm{i}\frac{1}{\sqrt{N}}\bm{u}\cdot\tilde{\bm{X}}}\psi e^{-\mathrm{i}\frac{1}{\sqrt{N}}\bm{u}\cdot\tilde{\bm{X}}},\quad\tilde{\phi}_{\bm{u}/\sqrt{N}}\coloneqq e^{\mathrm{i}\frac{1}{\sqrt{N}}\bm{u}\cdot\tilde{\bm{X}}^{\prime}}\phi e^{-\mathrm{i}\frac{1}{\sqrt{N}}\bm{u}\cdot\tilde{\bm{X}}^{\prime}}$. (297)
[…]
$(C)_{ki}$ $\coloneqq-\mathrm{i}\braket{k|X_{i}|\psi},\quad(C^{\prime})_{l^{\prime}i}\coloneqq-\mathrm{i}\braket{l^{\prime}|X_{i}^{\prime}|\phi},$ (298)
@@
where $C^{+}$ denotes the Moore–Penrose inverse of $C$. Note that the assumption that $\mathrm{Cov}(\psi,\bm{X})\geq\mathrm{Cov}(\phi,\bm{X}^{\prime})$ is equivalent to
@@
Let us define Hermitian operators $O,O^{\prime}$ by $O\coloneqq\bm{u}\cdot\tilde{\bm{X}}/\sqrt{N}$ and $O^{\prime}\coloneqq\bm{u}\cdot\tilde{\bm{X}}^{\prime}/\sqrt{N}$. Introducing a finite non-negative number $L$ by
$L\coloneqq\max\left\{\max_{\|\bm{\theta}\|=1}\|\bm{\theta}\cdot\tilde{\bm{X}}\|_{\infty},\max_{\|\bm{\theta}\|=1}\|\bm{\theta}\cdot\tilde{\bm{X}}^{\prime}\|_{\infty}\right\}$, (307)
we have $\|O\|_{\infty}\leq L\frac{\|\bm{u}\|}{\sqrt{N}}$ and $\|O^{\prime}\|_{\infty}\leq L\frac{\|\bm{u}\|}{\sqrt{N}}$. Applying Lemma 28 and Lemma 29 for $f(\theta)\coloneqq e^{-\mathrm{i}\theta O^{\prime}}\mathcal{E}(e^{\mathrm{i}\theta O}\psi e^{-\mathrm{i}\theta O})e^{\mathrm{i}\theta O^{\prime}}$, we get
@@
$\leq\frac{2^{3}}{6}(\|O\|_{\infty}+\|O^{\prime}\|_{\infty})^{3}\leq\frac{2^{6}}{6}\left(L\frac{\|\bm{u}\|}{\sqrt{N}}\right)^{3},$ (309)
@@
$\mathcal{E}(O\psi)$ $|,\displaystyle=\frac{-\mathrm{i}}{\sqrt{N}}\sum_{k^{\prime}=1}^{d^{\prime}-1}\left(ZC\bm{u}\right)_{k^{\prime}}\ket{k^{\prime}}\bra{\phi}$, (311)
@@
$\frac{1}{2}\braket{\phi|(O^{\prime 2}\mathcal{E}(\psi)+\mathcal{E}(\psi)O^{\prime 2})|\phi}$ $=\braket{\phi|O^{\prime 2}|\phi}=\frac{1}{N}\sum_{i,j=1}^{m}u^{i}u^{j}\braket{\phi|\tilde{X}^{\prime}_{i}\tilde{X}^{\prime}_{j}|\phi}=\frac{1}{N}(C^{\prime{\dagger}}\bm{u})^{\dagger}C^{\prime}\bm{u},$ (313)
@@
$\braket{\phi|O^{\prime}\mathcal{E}(O\psi)|\phi}$ $=\frac{1}{N}(C^{\prime}\bm{u})^{\dagger}ZC\bm{u},\quad\braket{\phi|\mathcal{E}(\psi O)O^{\prime}|\phi}=\frac{1}{N}(ZC\bm{u})^{\dagger}C^{\prime}\bm{u}.$ (317)
@@
$\frac{1}{2}\braket{\phi|\mathcal{E}(O^{2}\psi+\psi O^{2})|\phi}$ $=\frac{1}{N}(C\bm{u})^{\dagger}C\bm{u},\quad\braket{\phi|\mathcal{E}(O\psi O)|\phi}=\frac{1}{N}(C\bm{u})^{\dagger}\Gamma(C\bm{u}).$ (320)
@@
$=\frac{1}{N}\left(-(C^{\prime{\dagger}}\bm{u})^{\dagger}C^{\prime}\bm{u}+(C^{\prime}\bm{u})^{\dagger}ZC\bm{u}+(ZC\bm{u})^{\dagger}C^{\prime}\bm{u}-(C\bm{u})^{\dagger}C\bm{u}+(C\bm{u})^{\dagger}\Gamma(C\bm{u})\right)$ (322)
$=-\frac{1}{N}\bm{u}^{\dagger}(C^{\prime}-ZC)^{\dagger}(C^{\prime}-ZC)\bm{u}.$ (323)
@@
Since $\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{\bm{u}/\sqrt{N}}),\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)=\braket{\phi|f(1)|\phi}$, from Eqs. (309), (310), (312), and (325), we have proven
$\left|\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{\bm{u}/\sqrt{N}}),\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)-1\right|\leq\frac{2^{6}}{6}\left(L\frac{\|\bm{u}\|}{\sqrt{N}}\right)^{3}\leq\frac{2^{6}}{6}(LN^{-1/2+\epsilon})^{3}$ (326)
for any $\bm{u}$ such that $\|\bm{u}\|<N^{\epsilon}$.
Since $N^{-1/2+\epsilon}=o(1/N)$ if $\epsilon\in(0,1/6)$, we get
$\inf_{\|\bm{u}\|<N^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}^{\otimes N}(\tilde{\psi}_{\bm{u}/\sqrt{N}}^{\otimes N}),\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=\inf_{\|\bm{u}\|<N^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{\bm{u}/\sqrt{N}}),\tilde{\phi}_{\bm{u}/\sqrt{N}}\right)^{N}\geq\left(1+o\left(\frac{1}{N}\right)\right)^{N},$ (327)
@@
$\lim_{N\to\infty}\inf_{\|\bm{u}\|<N^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}^{\otimes N}(\tilde{\psi}_{\bm{u}/\sqrt{N}}^{\otimes N}),\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=1$ (328)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}(\tilde{\psi}_{\bm{u}/\sqrt{N}}^{\otimes N}),\tilde{\phi}_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=0.$ (329)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}(\psi_{\bm{u}/\sqrt{N}}^{\otimes N}),\phi_{\bm{u}/\sqrt{N}}^{\otimes N}\right)=0$ (330)
@@
For sets of Hermitian operators $\bm{X}\coloneqq\{X_{i}\}_{i=1}^{m}$ and $\bm{X}^{\prime}\coloneqq\{X_{i}^{\prime}\}_{i=1}^{m}$, we define $\mathcal{U}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}}$ and $\mathcal{U}^{\prime}_{\bm{\theta}}(\cdot)\coloneqq e^{\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}(\cdot)e^{-\mathrm{i}\bm{\theta}\cdot\bm{X}^{\prime}}$ for $\bm{\theta}\in\mathbb{R}^{m}$. For given two pure states $\psi$ and $\phi$, we define pure-state statistical models by $\mathcal{U}_{\bm{\theta}}(\psi)$ and $\mathcal{U}^{\prime}_{\bm{\theta}}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi}$ and $\mathcal{Q}^{\phi}$, respectively. If $r>0$ satisfies $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$, then there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{\bm{u}}{\sqrt{N}}}(\psi)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{\bm{u}}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}\right)=0$ (331)
@@
The assumption $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$ is equivalent to $\mathrm{Cov}(\psi,\bm{X})\geq r\,\mathrm{Cov}(\phi,\bm{X}^{\prime})$. Lemma 33 implies that there is a channel $\mathcal{E}$ such that
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\left(\mathcal{E}\left(e^{\mathrm{i}\frac{\bm{u}\cdot\bm{X}}{\sqrt{N}}}\psi e^{-\mathrm{i}\frac{\bm{u}\cdot\bm{X}}{\sqrt{N}}}\right)\right)^{\otimes N},\left(e^{\mathrm{i}\frac{\bm{u}\cdot\sqrt{r}\bm{X}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{\bm{u}\cdot\sqrt{r}\bm{X}^{\prime}}{\sqrt{N}}}\right)^{\otimes N}\right)=0$ (332)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}^{\prime}_{N}\left(e^{\mathrm{i}\frac{\bm{u}\cdot\sqrt{r}\bm{X}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{\bm{u}\cdot\sqrt{r}\bm{X}^{\prime}}{\sqrt{N}}}\right)^{\otimes N},\left(e^{\mathrm{i}\frac{\bm{u}\cdot\bm{X}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{\bm{u}\cdot\bm{X}^{\prime}}{\sqrt{N}}}\right)^{\otimes\lfloor rN\rfloor}\right)=0.$ (333)
@@
$\lim_{N\to\infty}\sup_{\|\bm{u}\|<N^{\epsilon}}T\left(\mathcal{E}^{\prime}_{N}\circ\mathcal{E}^{\otimes N}\left(e^{\mathrm{i}\frac{\bm{u}\cdot\bm{X}}{\sqrt{N}}}\psi e^{-\mathrm{i}\frac{\bm{u}\cdot\bm{X}}{\sqrt{N}}}\right)^{\otimes N},\left(e^{\mathrm{i}\frac{\bm{u}\cdot\bm{X}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{\bm{u}\cdot\bm{X}}{\sqrt{N}}}\right)^{\otimes\lfloor rN\rfloor}\right)=0.$ (334)
C.3 Proof of measurability of the set defined in Eq. (47)
@@
$H(\delta)\coloneqq\{\hat{g}\in G\mid\exists\bm{\theta},\,\mathcal{U}_{\bm{\theta}}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi),\,\|\bm{\theta}\|\leq\delta\}$ (335)
and prove that $H(\delta)$ is a closed set. Let $\{h_{n}\}_{n\in\mathbb{N}}$ be a convergent sequence in $H(\delta)$ such that $h_{n}\to h$ as $n\to\infty$. From the definition of $H(\delta)$, there exists a sequence $\{\bm{\theta}_{n}\}_{n\in\mathbb{N}}$ such that $\mathcal{U}_{\bm{\theta}_{n}}\circ\mathcal{U}_{h_{n}}(\psi)=\mathcal{U}_{g}(\psi)$ and $\|\bm{\theta}_{n}\|\leq\delta$. Due to the compactness of the closed ball of radius $\delta$, we can extract a convergent subsequence $\{\bm{\theta}_{n_{k}}\}_{k\in\mathbb{N}}$ such that $\bm{\theta}_{n_{k}}\to\bm{\theta_{0}}$ for some $\bm{\theta_{0}}$ as $k\to\infty$. Then, we have
$\mathcal{U}_{g}(\psi)=\mathcal{U}_{\bm{\theta}_{n_{k}}}\circ\mathcal{U}_{h_{n_{k}}}(\psi)\to\mathcal{U}_{\bm{\theta}_{0}}\circ\mathcal{U}_{h}(\psi),\qquad\delta\geq\|\bm{\theta}_{n_{k}}\|\to\|\bm{\theta}_{0}\|$ (336)

### v4

@@ D.2.2 Proof of Lemma 25
$(C)_{ki}$ $\coloneqq-\mathrm{i}\braket{k|X_{\mu}|\psi},\quad(C^{\prime})_{l^{\prime}i}\coloneqq-\mathrm{i}\braket{l^{\prime}|X_{\mu}^{\prime}|\phi},$ (300)
@@
where $C^{+}$ denotes the Moore–Penrose inverse of $C$. Note that the assumption that $\mathrm{Cov}(\psi,\{X_{\mu}\}_{\mu=1}^{m})\geq\mathrm{Cov}(\phi,\{X_{\mu}^{\prime}\}_{\mu=1}^{m})$ is equivalent to
@@
Let us define Hermitian operators $O,O^{\prime}$ by $O\coloneqq u^{\mu}\tilde{X}_{\mu}/\sqrt{N}$ and $O^{\prime}\coloneqq u^{\mu}\tilde{X}_{\mu}^{\prime}/\sqrt{N}$. Introducing a finite non-negative number $L$ by
$L\coloneqq\max\left\{\max_{\|\theta\|=1}\|\theta^{\mu}\tilde{X}_{\mu}\|_{\infty},\max_{\|\theta\|=1}\|\theta^{\mu}\tilde{X}_{\mu}^{\prime}\|_{\infty}\right\},$ (309)
we have $\|O\|_{\infty}\leq L\frac{\|u\|}{\sqrt{N}}$ and $\|O^{\prime}\|_{\infty}\leq L\frac{\|u\|}{\sqrt{N}}$. Applying Lemma 28 and Lemma 29 for $f(\theta)\coloneqq e^{-\mathrm{i}\theta O^{\prime}}\mathcal{E}(e^{\mathrm{i}\theta O}\psi e^{-\mathrm{i}\theta O})e^{\mathrm{i}\theta O^{\prime}}$, we get
@@
$\leq\frac{2^{3}}{6}(\|O\|_{\infty}+\|O^{\prime}\|_{\infty})^{3}\leq\frac{2^{6}}{6}\left(L\frac{\|u\|}{\sqrt{N}}\right)^{3},$ (311)
@@
$\mathcal{E}(O\psi)$ $|,\displaystyle=\frac{-\mathrm{i}}{\sqrt{N}}\sum_{k^{\prime}=1}^{d^{\prime}-1}\left(ZCu\right)_{k^{\prime}}\ket{k^{\prime}}\bra{\phi}$, (313)
@@
$\frac{1}{2}\braket{\phi|(O^{\prime 2}\mathcal{E}(\psi)+\mathcal{E}(\psi)O^{\prime 2})|\phi}$ $=\braket{\phi|O^{\prime 2}|\phi}=\frac{1}{N}u^{\mu}u^{\nu}\braket{\phi|\tilde{X}^{\prime}_{\mu}\tilde{X}^{\prime}_{\nu}|\phi}=\frac{1}{N}(C^{\prime{\dagger}}u)^{\dagger}C^{\prime}u,$ (315)
@@
$\braket{\phi|O^{\prime}\mathcal{E}(O\psi)|\phi}$ $=\frac{1}{N}(C^{\prime}u)^{\dagger}ZCu,\quad\braket{\phi|\mathcal{E}(\psi O)O^{\prime}|\phi}=\frac{1}{N}(ZCu)^{\dagger}C^{\prime}u.$ (319)
@@
$\frac{1}{2}\braket{\phi|\mathcal{E}(O^{2}\psi+\psi O^{2})|\phi}$ $=\frac{1}{N}(Cu)^{\dagger}Cu,\quad\braket{\phi|\mathcal{E}(O\psi O)|\phi}=\frac{1}{N}(Cu)^{\dagger}\Gamma(Cu).$ (322)
@@
$=\frac{1}{N}\left(-(C^{\prime{\dagger}}u)^{\dagger}C^{\prime}u+(C^{\prime}u)^{\dagger}ZCu+(ZCu)^{\dagger}C^{\prime}u-(Cu)^{\dagger}Cu+(Cu)^{\dagger}\Gamma(Cu)\right)$ (324)
$=-\frac{1}{N}u^{\dagger}(C^{\prime}-ZC)^{\dagger}(C^{\prime}-ZC)u.$ (325)
@@
Since $\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{u/\sqrt{N}}),\tilde{\phi}_{u/\sqrt{N}}\right)=\braket{\phi|f(1)|\phi}$, from Eqs. (311), (312), (314), and (327), we have proven
$\left|\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{u/\sqrt{N}}),\tilde{\phi}_{u/\sqrt{N}}\right)-1\right|\leq\frac{2^{6}}{6}\left(L\frac{\|u\|}{\sqrt{N}}\right)^{3}\leq\frac{2^{6}}{6}(LN^{-1/2+\epsilon})^{3}$ (328)
for any $u$ such that $\|u\|<N^{\epsilon}$.
Since $(N^{-1/2+\epsilon})^{3}=o(1/N)$ if $\epsilon\in(0,1/6)$, we get
$\inf_{\|u\|<N^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}^{\otimes N}(\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}),\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)=\inf_{\|u\|<N^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}(\tilde{\psi}_{u/\sqrt{N}}),\tilde{\phi}_{u/\sqrt{N}}\right)^{N}\geq\left(1+o\left(\frac{1}{N}\right)\right)^{N},$ (329)
@@
$\lim_{N\to\infty}\inf_{\|u\|<N^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}^{\otimes N}(\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}),\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)=1$ (330)
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}(\tilde{\psi}_{u/\sqrt{N}}^{\otimes N}),\tilde{\phi}_{u/\sqrt{N}}^{\otimes N}\right)=0.$ (331)
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}^{\otimes N}(\psi_{u/\sqrt{N}}^{\otimes N}),\phi_{u/\sqrt{N}}^{\otimes N}\right)=0$ (332)
@@
For sets of Hermitian operators $\{X_{\mu}\}_{\mu=1}^{m}$ and $\{X_{\mu}^{\prime}\}_{\mu=1}^{m}$, we define $\mathcal{U}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ and $\mathcal{U}^{\prime}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}^{\prime}}$ for $\theta\in\mathbb{R}^{m}$. For given two pure states $\psi$ and $\phi$, we define pure-state statistical models by $\mathcal{U}_{\theta}(\psi)$ and $\mathcal{U}^{\prime}_{\theta}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi}$ and $\mathcal{Q}^{\phi}$, respectively. If $r>0$ satisfies $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$, then there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}\right)=0$ (333)
@@
The assumption $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$ is equivalent to $\mathrm{Cov}(\psi,\{X_{\mu}\}_{\mu=1}^{m})\geq r\,\mathrm{Cov}(\phi,\{X_{\mu}^{\prime}\}_{\mu=1}^{m})$. Lemma 33 implies that there is a channel $\mathcal{E}$ such that
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\left(\mathcal{E}\left(e^{\mathrm{i}\frac{u^{\mu}X_{\mu}}{\sqrt{N}}}\psi e^{-\mathrm{i}\frac{u^{\mu}X_{\mu}}{\sqrt{N}}}\right)\right)^{\otimes N},\left(e^{\mathrm{i}\frac{\sqrt{r}u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{\sqrt{r}u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\right)^{\otimes N}\right)=0$ (334)
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}^{\prime}_{N}\left(e^{\mathrm{i}\frac{\sqrt{r}u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{\sqrt{r}u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\right)^{\otimes N},\left(e^{\mathrm{i}\frac{u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\right)^{\otimes\lfloor rN\rfloor}\right)=0.$ (335)
@@
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}^{\prime}_{N}\circ\mathcal{E}^{\otimes N}\left(e^{\mathrm{i}\frac{u^{\mu}X_{\mu}}{\sqrt{N}}}\psi e^{-\mathrm{i}\frac{u^{\mu}X_{\mu}}{\sqrt{N}}}\right)^{\otimes N},\left(e^{\mathrm{i}\frac{u^{\mu}X_{\mu}^{\prime}}{\sqrt{N}}}\phi e^{-\mathrm{i}\frac{u^{\mu}X_{\mu}}{\sqrt{N}}}\right)^{\otimes\lfloor rN\rfloor}\right)=0.$ (336)
D.2.3 Proof of measurability of the set defined in Eq. (231)
@@
$H(\delta)\coloneqq\{\hat{g}\in G\mid\exists\theta,\,\mathcal{U}_{\theta}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi),\,\|\theta\|\leq\delta\}$ (337)
and prove that $H(\delta)$ is a closed set. Let $\{h_{n}\}_{n\in\mathbb{N}}$ be a convergent sequence in $H(\delta)$ such that $h_{n}\to h$ as $n\to\infty$. From the definition of $H(\delta)$, there exists a sequence $\{\theta_{n}\}_{n\in\mathbb{N}}$ such that $\mathcal{U}_{\theta_{n}}\circ\mathcal{U}_{h_{n}}(\psi)=\mathcal{U}_{g}(\psi)$ and $\|\theta_{n}\|\leq\delta$. Due to the compactness of the closed ball of radius $\delta$, we can extract a convergent subsequence $\{\theta_{n_{k}}\}_{k\in\mathbb{N}}$ such that $\theta_{n_{k}}\to\theta_{0}$ for some $\theta_{0}$ as $k\to\infty$. Then, we have
$\mathcal{U}_{g}(\psi)=\mathcal{U}_{\theta_{n_{k}}}\circ\mathcal{U}_{h_{n_{k}}}(\psi)\to\mathcal{U}_{\theta_{0}}\circ\mathcal{U}_{h}(\psi),\qquad\delta\geq\|\theta_{n_{k}}\|\to\|\theta_{0}\|$ (338)